Practice problem six i'm lost on. Write the equation of the line passing through each of the given points with the indicated slope. Give results in slope-intercept form, if possible.

(0,5), m = - 3/5

Answer by tutorcecilia(2152)

This problem involves the equation of a line. The formula is y = mx + b
y = any value for y
m = the slope of the line
x = any matching value for x
b = the point where the line crosses the y-axis (the y-intercept)
.
Given that: (x, y)=(0,5) and m = - 3/5
.
So, plugging in the values:
y=5
m=-3/5
x=0
b=(solve for b)
.
y = m x + b
5=(-3/5)(0)+ b
5=b
.
Plug the value of the slope and the y-intercept(b) back into the formula:
Answer: y=-3/5x+5